OpenStudy (anonymous):

Solve for b 1 of a trapezoid in which A = (48x + 68) in 2 , h = 8 in., and b 2 = (9x + 12) in.

OpenStudy (anonymous):

Can anyone help me? 3.41 x 1,3019234 divided by 12347 x 8 + your age - 1804351 x 235-135425 squaredx 10 to the fifth power + 1- 1243818723847

OpenStudy (anonymous):

A= .5 h ( b1+ b2) 48x +68 = .5 (8)( b1 + 9x + 12) 48x + 68 = 4 ( b1 +9x +12) 48 x + 68 = 4 b1 + 36x + 48) 12x + 20 = 4 b1 3x + 5 = b1